# Mamba State-Space Models Can Be Strong Downstream Learners

### v1

@@ Abstract
Mamba [22] state-space models (SSMs) have recently outperformed state-of-the-art (SOTA) Transformer large language models (LLMs) in various tasks and been widely adapted. However, Mamba’s downstream learning capabilities remain either unexplored–e.g., mixed-precision (MPFT) and parameter-efficient fine-tuning (PEFT)–or under-evaluated–e.g., in-context learning (ICL). For the latter, recent works [42, 19] reported Mamba’s ICL rivals SOTA Transformer LLMs using non-standard benchmarks. In contrast, we show that on standard benchmarks, pretrained Mamba models achieve only 38% of the ICL performance improvements (over zero-shot) of comparable Transformers.
Enabling MPFT and PEFT in Mamba architectures is challenging due to recurrent dynamics and highly customized CUDA kernels, respectively. However, we prove that Mamba’s recurrent dynamics are robust to small input changes using dynamical systems theory. Empirically, we show that performance changes in Mamba’s inference and fine-tuning due to mixed-precision align with Transformer LLMs. Furthermore, we show that targeting key memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus achieving parameter efficiency while retaining speedups. We show that combining MPFT and PEFT enables up to 2.15 times more tokens-per-second and 65.5% reduced per-token-memory compared to full Mamba fine-tuning, while achieving up to 81.5% of the ICL performance improvements (over zero-shot) of comparably fine-tuned Transformers.
**footnotetext: Equal Contribution. Correspondence to john.t.halloran@leidos.com.
@@ 1 Introduction
Innovating on previous state-space models (SSMs) [23, 11], Mamba [22] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs). Mamba was initially shown to greatly outperform comparable Transformer LLMs [5] across a large number of standard natural language benchmarks. Subsequently, pretrained Mamba models have been widely adapted across different data modalities [39, 62, 34, 43, 35], tasks [57, 59, 45, 60, 54, 35, 2], and architectures [1, 42, 37].
However, despite such rapid and widespread adaptation, evaluation of Mamba’s ability to perform standard downstream learning abilities exhibited by Transformer-based LLMs have either not been extensively conducted on standard natural benchmarks or are completely lacking. For instance, while recent works [42, 19, 29] have evaluated Mamba’s ability to perform in-context learning (ICL), such studies focused extensively on either non-natural tasks [29, 17] or non-standard benchmarks [25].
Furthermore, evaluation of Mamba’s mixed-precision fine-tuning (MPFT) and performance efficient fine-tuning (PEFT) capabilities are currently lacking. For the former, MPFT (and, by extension, mixed-precision inference) are made difficult due to potential sensitivities of Mamba’s recurrent dynamics, where  [21, 28] suggest full precision (FP32) is required to perform stable training. For the latter, PEFT via standard low-rank adaptation (LoRA) [27] is made difficult within Mamba’s SSM layer (referred to herein as the MambaBlock) due to highly customized SSM CUDA kernels which provide competitive performance to attention-based speedups [10] at the cost of standard adapter support. However, PEFT and MPFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [52], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12].
Herein, we extensively explore Mamba’s downstream learning capabilities across standard natural benchmarks. For ICL, we show that, in contrast to recent non-standard studies showing Mamba models rival state-of-the-art (SOTA) LLMs of similar parameter counts, the pretrained benefits of Mamba few-shot learning are significantly less than comparable Transformer LLMs across standard natural benchmarks; averaged across the benchmarks and parameter counts in Table 1, Mamba models only achieve 38% of the performance improvements (relative to zero-shot) of comparable Transformer models from the Pythia suite [5]. However, we show in the sequel that Mamba models can more than halve this gap through efficient fine-tuning, achieving as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers.
@@
For MPFT, we leverage theory from dynamical systems to show that small input changes in a MambaBlock do not lead to exponentially deviating outputs. Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference and fine-tuning are on par with those demonstrated by Transformer LLMs (Section 7). For PEFT, we show that by targeting the largest memory buffer exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying. We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT.
Table 1: In-context learning performance for pretrained Mamba and Pythia models. Models are collected into parameter classes for head-to-head comparison using the groupings in [22]. Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16]. LAMBADA zero-shot is more effective for the model sizes considered (further discussed in [58, 8]) and thus excluded from few-shot performance averages. Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
Model $N$-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
Mamba 130M 0 16.07 44.3 35.3 64.5 48.0 24.2 44.8
1 19.34 38.3 35.2 64.3 47.1 23.5 51.3 -1.4
3 23.13 35.4 35.1 65.1 49.0 24.0 50.7 -0.2
5 24.38 36.2 34.8 64.9 49.2 23.9 50.5 -0.5
[…]
1 47.21 28.2 30.6 62.2 43.4 23.7 49.3 -0.4
3 63.70 24.7 30.5 61.9 44.8 22.9 51.3 0.1
5 66.30 25.3 30.4 62.6 43.4 23.1 50.8 -0.2
Mamba 370M 0 8.14 55.6 46.5 69.5 55.0 27.9 55.5
1 9.74 49.8 45.9 69.3 57.4 26.5 54.6 -0.8
3 10.89 48.5 46.2 69.6 58.7 28.5 53.6 1.0
5 11.36 48.5 46.2 69.4 58.3 28.0 56.0 1.3
[…]
1 12.26 47.1 40.5 68.0 53.8 25.6 52.4 1.8
3 14.39 43.2 40.9 67.9 55.1 26.9 54.0 4.2
5 14.62 44.1 40.8 68.1 54.6 26.6 53.4 3.5
Mamba 790M 0 6.01 61.7 55.1 72.1 61.2 29.6 56.0
1 7.06 56.2 54.5 72.5 63.3 30.1 56.9 1.4
3 8.05 54.8 54.2 72.2 63.4 31.6 57.2 2.4
5 8.83 53.4 54.6 72.5 64.6 32.1 57.5 3.4
[…]
1 8.99 51.8 47.3 70.7 57.1 28.2 53.4 1.0
3 10.48 48.2 47.5 71.2 59.2 28.0 54.3 2.2
5 10.86 48.4 47.3 71.4 58.7 28.4 53.1 1.9
Mamba 1.4B 0 5.04 65.0 59.1 74.2 65.5 32.9 58.6
1 5.83 60.6 58.20 74.7 64.5 33.0 61.2 -0.5
3 6.62 58.9 58.8 73.7 66.1 34.4 60.9 0.6
5 6.98 58.4 59.0 74.0 66.4 35.5 60.5 1.4
[…]
1 6.96 56.3 52.1 71.4 62.0 29.5 57.5 1.4
3 7.89 54.4 52.6 70.9 63.9 31.1 56.8 2.9
5 8.02 54.4 52.8 71.0 63.2 31.3 57.8 3.3
Mamba 2.8B 0 4.23 69.2 66.2 75.2 69.7 36.3 63.4
1 5.01 63.9 65.7 75.5 69.8 37.2 63.7 0.6
3 5.53 63.0 65.5 75.2 70.8 38.1 64.8 1.6
5 5.70 62.7 66.2 76.2 70.9 38.3 64.6 2.1
[…]
1 5.66 60.9 59.4 73.8 66.8 34.8 59.0 1.7
3 6.20 59.1 59.9 74.7 67.4 34.9 60.8 2.9
5 6.52 59.1 60.2 74.5 67.1 35.0 61.3 3.1
@@ 2 Background
2.1 Downstream learning for LLMs
Since the release of the Transformer architecture [51], attention-based LLMs have exhibited several downstream learning abilities–in particular, PEFT, MPFT, and ICL–which allow the rapid adaptation of foundation models towards specific applications. PEFT using adapters [24] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters. Arguably the most widely used such PEFT method is LoRA [27], which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
@@ 2.1 Downstream learning for LLMs
ICL provides an adaptable alternative to fine-tuning. Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest. Given sufficiently large models and pretraining data [8, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting. ICL thus avoids computational and storage costs associated with fine-tuning and PEFT adapter weights, respectively, and rather changes the prompt to immediately change the model for downstream learning. However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [8, 38].
2.2 State-space Models
Structured state-space sequence (S4) models [23, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency. Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned. Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations). However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures [22].
Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time. Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations [14]. To enable hardware efficiency with time-varying SSM parameters, [22] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states.
@@ 3 Mamba state-space models
For model dimension $d$ and maximum input sequence length $T$, the MambaBlock defines state-space parameters $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}$. The matrix $\bm{\Delta}_{t}$ controls the discrete step-size. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{T}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{T}\in\mathbb{R}{}^{d}$:
@@
3.1 Hardware-aware optimizations
As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [11] are no longer applicable. However, by diagonality, each dimension may be computed in parallel. Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [7]. [15] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated. Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{n\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on.
4 Stable dynamics in the MambaBlock
@@
The proof of Theorem 1 is available in Appendix A, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
4.1 Consequences for automatic mixed-precision
During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ fed into the SSM scan (which are passed through a SwiGLU), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
@@ 4.1 Consequences for automatic mixed-precision
5 Regularized fine-tuning using LoRA in the MambaBlock
Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer. E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these invidiual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer). Rather than adjusting Mamba’s low-level CUDA kernels themselves to integrate LoRA within the highly optimized prefix scan, we note that $\mathbf{W}$ can instead be directly targeted. Doing so, we have the following, where the proof is available in Appendix B.

### v2

# Mamba State-Space Models Are Lyapunov-Stable Learners

@@ Abstract
Mamba [21, 11] state-space models (SSMs) were recently shown to outperform state-of-the-art (SOTA) Transformer large language models (LLMs) across various tasks. Despite subsequent widespread adaptation, little work has focused on Mamba LLMs’ amenability for fine-tuning frameworks ubiquitously used for Transformer-based LLMs, e.g., mixed-precision fine-tuning (MPFT) and parameter-efficient fine-tuning (PEFT). For the former, it currently remains an open question whether Mamba’s recurrent dynamics are robust to small input changes, such as those encountered during MPFT. Using dynamical systems theory (in particular, Lyapunov exponents), we answer this question in the affirmative. We empirically validate this result through several experiments, showing that Mamba SSMs are significantly more stable to changes introduced by mixed-precision than comparable Transformers, even when both MPFT and PEFT are combined. For PEFT, we show how targeting specific memory buffers in Mamba’s customized CUDA kernels for low-rank adaptation regularizes SSM parameters, thus providing both parameter efficient learning and computational savings. Finally, with both MPFT and PEFT enabled, we explore the impact of instruction tuning Mamba SSMs for in-context learning (ICL) on natural language tasks. While pretrained Mamba and Mamba-2 models only achieve 38% and 82% (respectively) of the ICL improvements of comparable Transformer-based LLMs, we show that instruction tuning allows Mamba models to narrow this gap to 81% and Mamba-2 models to skyrocket over this gap to 132%.
**footnotetext: Correspondence to john.t.halloran@leidos.com.
@@ 1 Introduction
Innovating on previous state-space models (SSMs) [22, 10], Mamba [21, 11] has been recently proposed as an accurate, sub-quadratic alternative to Transformer large language models (LLMs). Upon their introduction in [21], Mamba SSMs were shown to greatly outperform comparable attention-based LLMs [4] across a large number of standard natural language benchmarks. Subsequently, pretrained Mamba models have been widely adapted across a large number of data modalities [36, 31, 43, 32], tasks [57, 53], and architectures  [1, 40, 34]. Despite such widespread adaptation and subsequent research threads [11, 40, 54], little work has been done to understand the amenability of Mamba SSMs for widely used fine-tuning frameworks, such as mixed-precision fine-tuning (MPFT) [38] and parameter-efficient fine-tuning (PEFT) [23, 25].
MPFT and PEFT are arguably two of the most widely utilized techniques for LLM alignment [50] and customization [51], and are typically combined to drastically decrease hardware demands needed to fine-tune modern LLMs [12]. However, direct application of MPFT for Mamba SSMs is made difficult due to potential sensitivities of Mamba’s state-space dynamics, a common concern for recurrent-based deep models [41]. To combat this, both [26] and [20] suggest full precision (FP32) may be required to perform stable training for Mamba models. Thus, it is currently an open question whether Mamba’s recurrent dynamics are stable in the presence of small input deviations, such as those introduced in MPFT.
To answer this question, we leverage theory from dynamical systems. Deriving and bounding the Lyapunov exponents for both Mamba and Mamba-2 models, we show that small input changes within the SSM layer of either model do not lead to exponentially deviating outputs. Empirically, we validate this theoretical result; compared to full-precision, deviations due to mixed-precision for Mamba inference are on par with those demonstrated by Transformer LLMs, while deviations due to MPFT are significantly more stable than those of comparable Transformers (Section 5). Furthermore, this trend continues when MPFT and PEFT are combined, where Mamba SSMs again produce significantly smaller deviations compared to comparable Transformer LLMs.
@@
For PEFT, we show that by targeting the large memory buffers exploited by Mamba’s highly customized CUDA kernels, LoRA may be used for extremely efficient fine-tuning, while simultaneously regularizing the majority of Mamba’s SSM parameters via weight tying. We show that this leads to extremely efficient PEFT, resulting in up to 2.15 times faster training and 65.5% reduced memory compared to the largest evaluated Mamba model without MPFT or PEFT. Furthermore, this allows even the largest (2.8 billion parameter) Mamba LLMs to be fine-tuned on a single GPU with as little as 24GB of onboard memory.
Finally, using both MPFT and PEFT, we complement existing studies [40, 30] by exploring the ICL capabilities of instruction-tuned Mamba and Mamba-2 models on natural language tasks. In particular, the ICL capabilities of both pretrained Mamba and Mamba-2 models lag behind those of comparable Transformer models from the Pythia suite [4]; averaged across five standard natural language benchmarks and foundation model sizes, Mamba and Mamba-2 models only achieve 38% and 82%, respectively, of the performance improvements (relative to zero-shot) of Pythia models. However, after instruction-tuning, Mamba models are able to achieve as much as 81.5% of the average few-shot learning improvement (relative to zero-shot) of comparable Transformers, while Mamba-2 models push this to 132% of the ICL improvements achieved by Pythia models. We note that, similar to Transformer foundation models [55], (post) instruction tuning ICL appears as an emergent abilities for Mamba and Mamba-2 SSMs, manifesting for models of size 370 million parameters and greater, while failing to manifest for Mamba and Mamba-2 models of fewer parameter counts.
We note that herein, when describing a particular foundation model or result, we use the term “Mamba model” to refer to one of the original models released in [21] and “Mamba-2 model” to refer to models released in [11]. While there are subtle architectural differences between these two SSMs, they share important similarities which allow our theoretical results to extend to both sets of models. In particular, Mamba and Mamba-2 models share the same state-space equations, support for SSM matrices, and design scheme of storing the majority of SSM parameters in a large memory buffer. Thus, we synonymously use the term MambaBlock to refer to the SSM layer of both Mamba and Mamba-2 models.
@@ 2 Background
Mixed-precision and parameter-efficient fine-tuning. In the current era of extremely large foundation models, both MPFT and PEFT have become ubiquitous tools for the rapid adaptation of Transformer-based LLMs towards specific applications. PEFT using adapters [23] allows a large pretrained model to be efficiently adapted for a particular downstream task by freezing the full model and training only a small number of extra parameters. Arguably the most widely used such PEFT method is LoRA [25], which injects trainable low-rank matrices into Transformer layers to approximate weight updates.
@@
State-space Models. Structured state-space sequence (S4) models [22, 14] are SSMs which leverage linear time-invariant (LTI) systems to combine the computational advantages of Transformers–i.e., highly parallelizable training–and recurrent neural networks (RNNs)–i.e., subquadratic autoregressive inference using recurrency. Within the S4 layer, an input signal is discretized and LTI parameters representing the input’s latent dynamics are learned. Owing to the S4 block’s latent dynamics being LTI, the S4 block’s output may be thus compactly represented as a single convolution between the input and an SSM convolution kernel (a matrix whose entries are products of LTI learnable parameters resulting from unrolling the state-space equations). However, despite hardware efficiency and long-dependency-modeling improvements, LTI-based S4 models remained inferior to Transformers of comparable parameter-sizes for natural language tasks, even when augmenting S4 layers with attention-layers for hybrid architectures [21].
Innovating on these previous S4 approaches, Mamba utilizes time-varying parameters to model latent dynamics, thus broadening the ability to capture nuanced changes evolving in discrete-time. Without LTI dynamics, however, the input-output representation via the SSM convolution kernel is no longer applicable, thus voiding previous hardware-aware S4 optimizations [14]. To enable hardware efficiency with time-varying SSM parameters, [21] thus introduced extensively customized CUDA kernels which implement highly parallelized prefix sums to compute recurrent states. Subsequently, [11] considered the unrolled state-space equations and leveraged tensor contractions (i.e., einsum notation [45]) to efficiently calculate Mamba variables. The resulting Mamba-2 foundation models contained significantly larger latent-variable dimensions than the Mamba models of [21], while maintaining efficiency on modern GPU accelerators.
In-context learning. ICL provides an adaptable alternative to fine-tuning. Rather than fine-tune the LLM directly, ICL augments a prompt with $n$ relevant examples (called shots) preceding the query of interest. Given sufficiently large models and pretraining data [7, 55], Transformer LLMs have proven adept at learning new concepts on the fly provided such few-shot prompting. However, it is worth noting that ICL inference time increases dramatically as the number of shots grows (due to self-attention’s quadratic complexity) and PEFT (when possible) is known to produce more accurate downstream learning results [7, 35].
@@ 3 Mamba state-space models
For latent-variable dimension $d$ and maximum input sequence length $T$, the MambaBlock defines state-space parameters $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}\in\mathbb{R}{}^{d\times d}$ for $t\in\{1,\dots,T\}$. The matrix $\bm{\Delta}_{t}$ controls the discrete step-size. Given an input sequence $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in\mathbb{R}{}^{d}$, the following linear mapping through latent states $\bm{x}_{1},\dots,\bm{x}_{T}\in\mathbb{R}{}^{d}$ is used to produce the output $\mathbf{y}_{1},\dots,\mathbf{y}_{T}\in\mathbb{R}{}^{d}$:
@@
3.1 Stable dynamics in the MambaBlock
@@
The proof of Theorem 1 is available in Appendix B, where the maximal Lyapunov exponent for an arbitrary MambaBlock is first proven to be non-positive. The main result subsequently follows.
Thus, the latent states of Mamba and Mamba-2 models are stable under small input changes. However, variables $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ are the primary outputs for such models, particularly for LLM applications. We next show that, given Theorem 1, Mamba and Mamba-2 output variables are also stable.
Theorem 2.
Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. Then small changes to the output $\mathbf{y}_{t}$ are also exponentially non-increasing over discrete time.
The proof of Theorem 2 is available in Appendix C. Thus, by Theorems 1 and  2, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during recurrency.
3.1.1 Consequences for automatic mixed-precision
During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. In contrast to full-precision fine-tuning, MPFT within the MambaBlock thus results in small differences to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish), $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), and the gradients calculated during training.
@@ 3.1.1 Consequences for automatic mixed-precision
3.2 Hardware-aware optimizations and PEFT
As matrices $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ are time-varying, S4 optimizations via the SSM convolution kernel [10] are no longer applicable. However, by diagonality, each dimension may be computed in parallel. Furthermore, the recurrence along every dimension is a prefix sum (also called a scan), which is highly parallelizable [6]. [21] thus capitalizes on this through extensively customized CUDA kernels wherein the majority of temporal variables are carefully laid out in a large buffer of GPU memory and manipulated. Instantiated as a PyTorch linear layer’s weight matrix, this memory buffer $\mathbf{W}\in\mathbb{R}{}^{T\times 3d}$ is used to store and access the diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ for all $t\in\{1,\dots,T\}$, such that
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. The customized Mamba prefix scan kernel heavily relies on this memory layout to optimize the access pattern of $\mathbf{W}$ in Equations 1 and 2.
Similarly, Mamba-2 stores diagonal elements of $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$ in a large memory buffer $\mathbf{W}$. However, rather than utilizing the underlying recurrence to directly compute the hidden state and output at each time-step, [11] consider the matrix resulting from unrolling Equation 2 across all $t$. Mamba-2 models thus leverage the structure of the ensuing semiseparable matrix to calculate $\bm{x}_{1},\dots,\bm{x}_{T}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ using tensor contractions, which are highly optimized on modern hardware accelerators.
When fine-tuning using LoRA, low-rank matrices are used to adapt the frozen weight matrices of targeted linear layers. The importance of $\mathbf{W}$ for both Mamba and Mamba-2 models makes it a primary candidate for LoRA adaptation. In such cases, selecting $\mathbf{W}$ for LoRA adaptation results in the following: